Hello, please EXPLAIN how you got your answer. Thank you!

int x = 1

int j;

for (j=0; j <=2  ; j++)

x = x * j ;

cout << x << endl;

What is the output of the C++ code above?

Answer:

A) 0

B) 1

C) 2

D) 3
Answer #1
int x = 1;  // x is initialized to 1
int j;
for (j=0; j <=2; j++)   // here this for loop iterates 3 times. for j=0, j=1,j=2
    x = x * j;
// let's see what happens with the above for loop
// j=0
// x = x*j = x*0 = 0
// j=1
// x = x*j = 0*1 = 0
// j=2
// x = x*j = 0*2 = 0
// so, after the loop, value of x is 0
cout << x << endl;  // prints 0

Output:  0